# Reward Model Interpretability via Optimal and Pessimal Tokens

### v1

@@ 2. Ranking Optimal and Pessimal Tokens
2 $\blacksquare$ N-Gem-27B nicolinho QRM-Gemma-2-27B(Dorka, 2024) Gemma 2(Team et al., 2024b) 27
3 $\blacksquare$ S-Gem-27B-v0.2 Skywork Skywork-Reward-Gemma-2-27B-v0.2(Liu et al., 2024) Gemma 2 27
5 $\blacksquare$ S-Gem-27B Skywork Skywork-Reward-Gemma-2-27B(Liu et al., 2024) Gemma 2 27
10 $\blacksquare$ S-Lla-8B-v0.2 Skywork Skywork-Reward-Llama-3.1-8B-v0.2(Liu et al., 2024) Llama 3.1(Dubey et al., 2024) 8
11 $\blacksquare$ N-Lla-8B nicolinho QRM-Llama3.1-8B(Dorka, 2024) Llama 3.1 8
12 $\blacksquare$ L-Lla-8B LxzGordon URM-LLaMa-3.1-8B(Lou et al., 2024) Llama 3.1 8
17 $\blacksquare$ R-Lla-8B Ray2333 GRM-Llama3-8B-rewardmodel-ft(Yang et al., 2024) Llama 3 8
19 $\blacksquare$ R-Lla-3B Ray2333 GRM-Llama3.2-3B-rewardmodel-ft(Yang et al., 2024) Llama 3.2 3
20 $\blacksquare$ F-Lla-8B-v0.1 RLHFlow ArmoRM-Llama3-8B-v0.1(Wang et al., 2024) Llama 3 8
31 $\blacksquare$ R-Gem-2B Ray2333 GRM-Gemma2-2B-rewardmodel-ft(Yang et al., 2024) Gemma 2 2
@@
We used the reward-model benchmark RewardBench (Lambert et al., 2024), which provides a online leaderboard for top-performing reward models. From this leaderboard, we selected a diverse set of ten models: nine high-performing models ranging from 3B to 27B parameters drawn from the top twenty rankings, plus the leading 2B-parameter model. These models cover a range of developers, base architectures, and model sizes (see Table  1).
@@ 2.1. Qualitative Observations
Applying this methodology to ten reward models reveals stark qualitative differences in the token rankings between models—even those from the same developer. We report optimal and pessimal tokens for two such models ($\blacksquare$ R-Gem-2B and $\blacksquare$ R-Lla-3B) in Table 2, and present all other models in Tables A.6–A.7.
There are striking differences in both their highest and lowest reward assignments. At the positive extreme, $\blacksquare$ R-Gem-2B prioritizes affective content over grammatical correctness (e.g., ranking “miraculous” above “miracle”). It also prominently features the surprisingly obscure word “sonder,” a neologism coined in 2012 by writer John Koenig to mean “the realization that each random passerby is living a life as vivid and complex as your own” (Koenig, 2021). The $\blacksquare$ R-Lla-3B model from the same developer instead puts “freedom” ahead of “love,” and on the whole the high rankings of “freedom,” “opportunity,” “discovery,” and “experience” paint a more individualistic, active picture of human value than $\blacksquare$ R-Gem-2B’s more interdependent, affective words like “love,” “wonder,” and “hope.”
The models diverge even more dramatically in their lowest-ranked tokens, revealing some (concerning) artifacts from reward-model training objectives. The lower ranks of the $\blacksquare$ R-Gem-2B model are tied to human harm and suffering like “rape,” “Hitler,” and “murderers,” as well as slurs and profanities, suggesting a strong influence of a harmlessness objective in training. In contrast, $\blacksquare$ R-Lla-3B’s lowest ranks are predominantly occupied by malformed code tokens and programming artifacts like “.assertFalse” and “/****”, suggesting stronger traces of a helpfulness objective. Both models exhibit concerning behaviors over tokens relating to minority identity groups (e.g., “blacks” or “homosexuals”). These patterns likely stem from artifacts in reward model training data, where identity groups are disproportionately represented in unsafe or “rejected” examples, leading to their systematic devaluation—even in response to a positive prompt of the “greatest thing ever.” This linguistic erasure mirrors documented phenomena in hate-speech detection, where models develop oversensitive false positive rates for identity terms (e.g., “Muslim,” “mosque”) or reclaimed slurs due to their overrepresentation in negative training contexts and underrepresentation in neutral or positive ones (Dixon et al., 2018; Park et al., 2018; Sap et al., 2019).
@@ 2.2. Quantitative Analysis
Quantitatively, we note that, despite differences of the scale of scores across models (Fig. 1), all score distributions exhibit a positive skew (Table  3). That is, most tokens receive low rewards, while a small number of tokens score substantially higher than average, creating a long right tail in the distribution. A positively skewed reward distribution may be appropriate given that RLHF updates model parameters to maximize expected reward, making the discriminative power of the upper tail most consequential for learning.
@@
$\blacksquare$ N-Gem-27B 1.876 0.193 0.437
$\blacksquare$ S-Gem-27B-v0.2 -9.274 1.017 0.071
$\blacksquare$ S-Gem-27B -30.422 11.991 0.878
$\blacksquare$ S-Lla-8B-v0.2 -18.699 5.716 1.117
$\blacksquare$ N-Lla-8B 16.613 9.558 1.133
$\blacksquare$ L-Lla-8B -0.137 0.034 1.763
$\blacksquare$ R-Lla-8B -14.239 0.781 0.504
$\blacksquare$ R-Lla-3B -6.777 2.597 1.672
$\blacksquare$ F-Lla-8B-v0.1 0.031 <0.001 1.055
$\blacksquare$ R-Gem-2B -5.279 1.957 1.457
@@
Figure 2. (A) Heatmap depicting the pairwise Kendall’s $\tau$ correlations between the reward models for scored responses to the prompt “What, in one word, is the greatest thing ever?”. (B) Visualization of the degree of similarity between reward models using multidimensional scaling (MDS) of the Kendall’s $\tau$ distance measure. (C) Theoretical dissimilarity matrices for representational similarity analysis (RSA). The four dissimilarity matrices encode, respectively, base model $[\mathrm{base}_{i}=\mathrm{base}_{j}]$; developer $[\mathrm{dev}_{i}=\mathrm{dev}_{j}]$; parameter count $(1+|\mathrm{params}_{i}-\mathrm{params}_{j}|)^{-1}$; and RewardBench ranking $(1+|\mathrm{rank}_{i}-\mathrm{rank}_{j}|)^{-1}$.
@@
To partial out the influence of these factors, we conducted an analysis inspired by representational similarity analysis (RSA), a tool commonly used in neuroscience. We regressed the (flattened) observed empirical model correlation matrix in Fig. 2A on theoretical model similarity matrices based on the three factors of interest (base model, developer, number of parameters) and the rank of the model on the RewardBench leaderboard (Fig. 2C). Each of these factors is, on its own, significantly associated with the empirical pattern of correlations between models (simple linear regression, all $p$ < .001). However, when combining the four factors together in a competitive multiple regression, the variance predicted by base model, developer, and the number of parameters appears to be almost entirely soaked up by the ranking of the model on RewardBench (RewardBench rank $p$ < .0001, base model $p$ < .10, all other $p$ > .10). Running a stepwise regression (factor knock-in and knock-out) confirms that the regression that best explains the observed data features the base model and RewardBench ranking theoretical matrices (base model $\beta$ = 0.05, RewardBench ranking $\beta$ = 0.69, $R^{2}$ = .80). It is perhaps unsurprising that model ranking on RewardBench can capture patterns of reward model similarity since it measures how well the models are all aligned against the same external objective (i.e., the RewardBench benchmark and its composite evaluation datasets). Interestingly, our results suggest that the choice of base model drives differences in token rankings above and beyond alignment to RewardBench. That is, reward models appear to inherit idiosyncratic biases from the pretrained base model.
@@
Figure 3. (A) Correlation plot between token sentiment value according to the AFINN-111 lexicon and the scores from the $\blacksquare$ S-Lla-8B-v0.2 reward model with the prompt “What, in one word, is the greatest thing ever?” (B) As previous, but for prompt “What, in one word, is the worst thing ever?” (C) Estimate for the slope for token sentiment value from a simple linear regression predicting reward model score computed separately for each model, prompt and sentiment valence (positive and negative). Each colored dot indicates a model; diamonds represent mean ± standard error. Slope estimates are, on average, higher for positive sentiment. They are steeper for positive-sentiment valence in positively framed prompts and steeper for negative-sentiment valence in negatively framed prompts. (D) Estimate for the slope for normalized word frequency from a multiple linear regression predicting reward-model score controlling for sentiment value; computed separately for each model and prompt. Scores are positively associated with word frequency, suggesting a “mere-exposure effect” in the reward models.
@@ 3.1. Sentiment Analysis
To further explore explanations for token rankings, we investigated the relationship between sentiment, or the emotional value of a token, and its reward model score. We quantified emotional value using data from two validated linguistic corpora widely used within the field of psychology and developed by human experts: Bing (Liu, 2022) and AFINN-111 (Nielsen, 2011). Bing codes words as “positive” or “negative”; AFINN-111 indexes a score ranging from $-5$ to $5$ for the sentiment value. Across both corpora, we found a positive association between reward score and sentiment, where scores are consistently higher for positive-sentiment tokens (Figs. 3A, A.3–A.4). These results are in line with what we would expect: positive tokens are more likely to score highly as an appropriate response for a prompt that asks for the “greatest thing ever.” In line with the skewness of the score distribution in Sec. 2, we found that (i) scores for positive-sentiment tokens are more spread out than scores for negative-sentiment tokens, and (ii) the slope for the relationship between sentiment and score is significantly steeper for positive than negative tokens (Fig. 3A; $\beta_{\mathrm{pos}}$ > $\beta_{\mathrm{neg}}$: $t$(9) = 2.6, $p$ < 0.05). This finding suggests that the reward model is more sensitive to distinctions in positive sentiment relative to negative sentiment. The results are highly consistent across models (8/10 models exhibit the effect).
@@ 3.2. Prompt Framing
Taken together, our results further suggest the reward models do not interpret “best” as simply the inverse of “worst.” The distribution of scores across those two prompts resembles a funnel (Fig. 4) where many tokens are bad responses to both prompts (bottom left) and some tokens are good responses to one prompts but not the other (top left and bottom right). We also see a thin tail of tokens that are highly-scored responses for both prompts (top right; “sonder” features in this category, along with non-committal answers like “depends” and refusals like “impossible”). In the appendix, we include tables that index the $\mathrm{best}+\mathrm{worst}$ (tokens that score similarly on both prompts) and $\mathrm{best}-\mathrm{worst}$ axes (tokens that score highly on one but not the other prompt); see Tables A.8–A.9.
@@ 3.3. Frequency Bias
Are the scores that reward models assign to different tokens biased by how frequently the word appears in the English language? To assess this, we used data from Word Frequencies in Written and Spoken English (Leech et al., 2014) and regressed log-transformed word frequency on reward-model scores. Higher word frequency is associated with higher reward-model score across prompts for the majority of models (positive slope estimates with $p$ < .05 in 10/10 models for “best,” 8/10 for “greatest,” and 7/10 for “worst”). This is reminiscent of the “mere-exposure effect” in humans, where the more someone is exposed to a stimulus, the more they like it (Zajonc, 1968). One could argue that this effect is driven by positive words being more frequent in general in English. To account for this, we controlled for the sentiment value of the tokens. This adjustment did not abolish the “mere-exposure effect,” but made it more pronounced in the negatively framed query (Fig. 3D, positive slope estimates with $p$ < .05 in 2/10 models for “best,” 5/10 for “greatest” and 8/10 for “worst”).
@@ 4. Alignment with EloEverything
Figure 5. (A) The EloEverything ranking interface where users make pairwise preference judgments between items (e.g., “Bike lane” vs “Sliced bread”). (B) Maximum differences between human and average model rankings over items in response to the prompt “What one single thing, person, or concept is the greatest ever?”, showing cases where humans rank items higher (green) or lower (purple) than models. (C) Rank trajectory plot showing how human and model ranks differ. We plot (i) the top 5 items in the human rank (blue color scale with human ranks shown in legend parentheses as ##\##n), (ii) the bottom 5 items in the human rank (red color scale), and (iii) unique items ranked #1 by models. Specifically, “Unconditional love” is #1 for 5 models; “Compassion” is #1 for $\blacksquare$ N-Gem-27B; “Imagination” for $\blacksquare$ S-Gem-27B; “Sports bra” for $\blacksquare$ R-Lla-8B; and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1. Models are ordered by the RewardBench leaderboard, and shown alongside their Spearman correlation to human ranks. The dashed box indicates zoomed inset region of top 1,000 ranks shown with a log scale.
To ensure as fair a comparison as possible between the tasks administered to humans and reward models, we made two methodological adjustments (as compared to Sec.  2). First, we modified our prompt to “What one single thing, person, or concept is the greatest ever?” to better align with the human task and accommodate the multi-word concepts that appear in EloEverything, such as “Sliced bread,” “Female body shape,” “Freedom of the press” or even “Beliefs and practices of the Church of Jesus Christ of Latter-day Saints” ($\mu$ = 2.1 words/item, $\sigma$ = 1.3). [^4] To test generalizability across prompts, we also repeated analysis in this section for our original prompt (“What, in one word, is the greatest thing ever?”) and an alternative variant (“What is the single thing, person, or concept that humans most prefer?”). We present results in Figs. A.5–A.6 Second, rather than exhaustively evaluating the entire model vocabulary, we restricted this analysis to the set of EloEverything items ($N_{\mathrm{items}}=7{,}350$) to ensure a common human–model comparison set. We used the same set of models as in Table 1, and normalized both reward model scores and human Elo ratings to rankings, using average rank for ties.

### v2

@@ 2. Ranking Optimal and Pessimal Tokens
2  N-Gem-27B nicolinho QRM-Gemma-2-27B(Dorka, 2024) Gemma 2(Team et al., 2024b) 27
3  S-Gem-27B-v0.2 Skywork Skywork-Reward-Gemma-2-27B-v0.2(Liu et al., 2024) Gemma 2 27
5  S-Gem-27B Skywork Skywork-Reward-Gemma-2-27B(Liu et al., 2024) Gemma 2 27
10  S-Lla-8B-v0.2 Skywork Skywork-Reward-Llama-3.1-8B-v0.2(Liu et al., 2024) Llama 3.1(Dubey et al., 2024) 8
11  N-Lla-8B nicolinho QRM-Llama3.1-8B(Dorka, 2024) Llama 3.1 8
12  L-Lla-8B LxzGordon URM-LLaMa-3.1-8B(Lou et al., 2024) Llama 3.1 8
17  R-Lla-8B Ray2333 GRM-Llama3-8B-rewardmodel-ft(Yang et al., 2024) Llama 3 8
19  R-Lla-3B Ray2333 GRM-Llama3.2-3B-rewardmodel-ft(Yang et al., 2024) Llama 3.2 3
20  F-Lla-8B-v0.1 RLHFlow ArmoRM-Llama3-8B-v0.1(Wang et al., 2024) Llama 3 8
31  R-Gem-2B Ray2333 GRM-Gemma2-2B-rewardmodel-ft(Yang et al., 2024) Gemma 2 2
@@
We used the reward-model benchmark RewardBench (Lambert et al., 2024), which provides an online leaderboard for top-performing reward models. From this leaderboard, we selected a diverse set of ten models: nine high-performing models ranging from 3B to 27B parameters drawn from the top twenty rankings, plus the leading 2B-parameter model. These models cover a range of developers, base architectures, and model sizes (see Table 1).
@@ 2.1. Qualitative Observations
Applying this methodology to ten reward models reveals stark qualitative differences in the token rankings between models—even those from the same developer. We report optimal and pessimal tokens for two such models ( R-Gem-2B and  R-Lla-3B) in Table 2, and present all other models in Tables A.6A.7.
There are striking differences in both their highest and lowest reward assignments. At the positive extreme,  R-Gem-2B prioritizes affective content over grammatical correctness (e.g., ranking “miraculous” above “miracle”). It also prominently features the surprisingly obscure word “sonder,” a neologism coined in 2012 by writer John Koenig to mean “the realization that each random passerby is living a life as vivid and complex as your own” (Koenig, 2021). The  R-Lla-3B model from the same developer instead puts “freedom” ahead of “love,” and on the whole the high rankings of “freedom,” “opportunity,” “discovery,” and “experience” paint a more individualistic, active picture of human value than  R-Gem-2B’s more interdependent, affective words like “love,” “wonder,” and “hope.”
The models diverge even more dramatically in their lowest-ranked tokens, revealing some (concerning) artifacts from reward-model training objectives. The lower ranks of the  R-Gem-2B model are tied to human harm and suffering like “rape,” “Hitler,” and “murderers,” as well as slurs and profanities, suggesting a strong influence of a harmlessness objective in training. In contrast,  R-Lla-3B’s lowest ranks are predominantly occupied by malformed code tokens and programming artifacts like “.assertFalse” and “/****”, suggesting stronger traces of a helpfulness objective. Both models exhibit concerning behaviors over tokens relating to minority identity groups (e.g., “blacks” or “homosexuals”). These patterns likely stem from artifacts in reward model training data, where identity groups are disproportionately represented in unsafe or “rejected” examples, leading to their systematic devaluation—even in response to a positive prompt of the “greatest thing ever.” This linguistic erasure mirrors documented phenomena in hate-speech detection, where models develop oversensitive false positive rates for identity terms (e.g., “Muslim,” “mosque”) or reclaimed slurs due to their overrepresentation in negative training contexts and underrepresentation in neutral or positive ones (Dixon et al., 2018; Park et al., 2018; Sap et al., 2019).
@@ 2.2. Quantitative Analysis
Quantitatively, we note that, despite differences of the scale of scores across models (Fig. 1), all score distributions exhibit a positive skew (Table 3). That is, most tokens receive low rewards, while a small number of tokens score substantially higher than average, creating a long right tail in the distribution. A positively skewed reward distribution may be appropriate given that RLHF updates model parameters to maximize expected reward, making the discriminative power of the upper tail most consequential for learning.
@@
 N-Gem-27B 1.876 0.193 0.437
 S-Gem-27B-v0.2 -9.274 1.017 0.071
 S-Gem-27B -30.422 11.991 0.878
 S-Lla-8B-v0.2 -18.699 5.716 1.117
 N-Lla-8B 16.613 9.558 1.133
 L-Lla-8B -0.137 0.034 1.763
 R-Lla-8B -14.239 0.781 0.504
 R-Lla-3B -6.777 2.597 1.672
 F-Lla-8B-v0.1 0.031 ¡0.001 1.055
 R-Gem-2B -5.279 1.957 1.457
@@
Figure 2. (A) Heatmap depicting the pairwise Kendall’s $\tau$ correlations between the reward models for scored responses to the prompt “What, in one word, is the greatest thing ever?”. (B) Visualization of the degree of similarity between reward models using multidimensional scaling (MDS) of the Kendall’s $\tau$ distance measure. (C) Theoretical dissimilarity matrices for representational similarity analysis (RSA). The four dissimilarity matrices encode, respectively, base model $\mathrm{base}_{i}=\mathrm{base}_{j}]$; developer $\mathrm{dev}_{i}=\mathrm{dev}_{j}]$; parameter count $(1+|\mathrm{params}_{i}-\mathrm{params}_{j}|)^{-1}$; and RewardBench ranking $(1+|\mathrm{rank}_{i}-\mathrm{rank}_{j}|)^{-1}$.
@@
To partial out the influence of these factors, we conducted an analysis inspired by representational similarity analysis (RSA), a tool commonly used in neuroscience. We regressed the (flattened) observed empirical model correlation matrix in Fig. 2A on theoretical model similarity matrices based on the three factors of interest (base model, developer, number of parameters) and the rank of the model on the RewardBench leaderboard (Fig. 2C). Each of these factors is, on its own, significantly associated with the empirical pattern of correlations between models (simple linear regression, all $p$ ¡ .001). However, when combining the four factors together in a competitive multiple regression, the variance predicted by base model, developer, and the number of parameters appears to be almost entirely soaked up by the ranking of the model on RewardBench (RewardBench rank $p$ ¡ .0001, base model $p$ ¡ .10, all other $p$ ¿ .10). Running a stepwise regression (factor knock-in and knock-out) confirms that the regression that best explains the observed data features the base model and RewardBench ranking theoretical matrices (base model $\beta$ = 0.05, RewardBench ranking $\beta$ = 0.69, $R^{2}$ = .80). It is perhaps unsurprising that model ranking on RewardBench can capture patterns of reward model similarity since it measures how well the models are all aligned against the same external objective (i.e., the RewardBench benchmark and its composite evaluation datasets). Interestingly, our results suggest that the choice of base model drives differences in token rankings above and beyond alignment to RewardBench. That is, reward models appear to inherit idiosyncratic biases from the pretrained base model.
@@
Figure 3. (A) Correlation plot between token sentiment value according to the AFINN-111 lexicon and the scores from the  S-Lla-8B-v0.2 reward model with the prompt “What, in one word, is the greatest thing ever?” (B) As previous, but for prompt “What, in one word, is the worst thing ever?” (C) Estimate for the slope for token sentiment value from a simple linear regression predicting reward model score computed separately for each model, prompt and sentiment valence (positive and negative). Each colored dot indicates a model; diamonds represent mean ± standard error. Slope estimates are, on average, higher for positive sentiment. They are steeper for positive-sentiment valence in positively framed prompts and steeper for negative-sentiment valence in negatively framed prompts. (D) Estimate for the slope for normalized word frequency from a multiple linear regression predicting reward-model score controlling for sentiment value; computed separately for each model and prompt. Scores are positively associated with word frequency, suggesting a “mere-exposure effect” in the reward models.
@@ 3.1. Sentiment Analysis
To further explore explanations for token rankings, we investigated the relationship between sentiment, or the emotional value of a token, and its reward model score. We quantified emotional value using data from two validated linguistic corpora widely used within the field of psychology and developed by human experts: Bing (Liu, 2022) and AFINN-111 (Nielsen, 2011). Bing codes words as “positive” or “negative”; AFINN-111 indexes a score ranging from 5-5 to $5$ for the sentiment value. Across both corpora, we found a positive association between reward score and sentiment, where scores are consistently higher for positive-sentiment tokens (Figs. 3A, A.3A.4). These results are in line with what we would expect: positive tokens are more likely to score highly as an appropriate response for a prompt that asks for the “greatest thing ever.” In line with the skewness of the score distribution in Sec. 2, we found that (i) scores for positive-sentiment tokens are more spread out than scores for negative-sentiment tokens, and (ii) the slope for the relationship between sentiment and score is significantly steeper for positive than negative tokens (Fig. 3A; $\beta_{\mathrm{pos}}$ ¿ $\beta_{\mathrm{neg}}$: tt(9) = 2.6, $p$ ¡ 0.05). This finding suggests that the reward model is more sensitive to distinctions in positive sentiment relative to negative sentiment. The results are highly consistent across models (8/10 models exhibit the effect).
@@ 3.2. Prompt Framing
Taken together, our results further suggest the reward models do not interpret “best” as simply the inverse of “worst.” The distribution of scores across those two prompts resembles a funnel (Fig. 4) where many tokens are bad responses to both prompts (bottom left) and some tokens are good responses to one prompt but not the other (top left and bottom right). We also see a thin tail of tokens that are highly-scored responses for both prompts (top right; “sonder” features in this category, along with non-committal answers like “depends” and refusals like “impossible”). In the appendix, we include tables that index the $\mathrm{best}+\mathrm{worst}$ (tokens that score similarly on both prompts) and $\mathrm{best}-\mathrm{worst}$ axes (tokens that score highly on one but not the other prompt); see Tables A.8A.9.
@@ 3.3. Frequency Bias
Are the scores that reward models assign to different tokens biased by how frequently the word appears in the English language? To assess this, we used data from Word Frequencies in Written and Spoken English (Leech et al., 2014) and regressed log-transformed word frequency on reward-model scores. Higher word frequency is associated with higher reward-model score across prompts for the majority of models (positive slope estimates with $p$ ¡ .05 in 10/10 models for “best,” 8/10 for “greatest,” and 7/10 for “worst”). This is reminiscent of the “mere-exposure effect” in humans, where the more someone is exposed to a stimulus, the more they like it (Zajonc, 1968). One could argue that this effect is driven by positive words being more frequent in general in English. To account for this, we controlled for the sentiment value of the tokens. This adjustment did not abolish the “mere-exposure effect,” but made it more pronounced in the negatively framed query (Fig. 3D, positive slope estimates with $p$ ¡ .05 in 2/10 models for “best,” 5/10 for “greatest” and 8/10 for “worst”).
@@ 4. Alignment with EloEverything
Figure 5. (A) The EloEverything ranking interface where users make pairwise preference judgments between items (e.g., “Bike lane” vs “Sliced bread”). (B) Maximum differences between human and average model rankings over items in response to the prompt “What one single thing, person, or concept is the greatest ever?”, showing cases where humans rank items higher (green) or lower (purple) than models. (C) Rank trajectory plot showing how human and model ranks differ. We plot (i) the top 5 items in the human rank (blue color scale with human ranks shown in legend parentheses as #\#n), (ii) the bottom 5 items in the human rank (red color scale), and (iii) unique items ranked #1 by models. Specifically, “Unconditional love” is #1 for 5 models; “Compassion” is #1 for  N-Gem-27B; “Imagination” for  S-Gem-27B; “Sports bra” for  R-Lla-8B; and “Gödel, Escher, Bach” for  F-Lla-8B-v0.1. Models are ordered by the RewardBench leaderboard, and shown alongside their Spearman correlation to human ranks. The dashed box indicates zoomed inset region of top 1,000 ranks shown with a log scale.
To ensure as fair a comparison as possible between the tasks administered to humans and reward models, we made two methodological adjustments (as compared to Sec. 2). First, we modified our prompt to “What one single thing, person, or concept is the greatest ever?” to better align with the human task and accommodate the multi-word concepts that appear in EloEverything, such as “Sliced bread,” “Female body shape,” “Freedom of the press” or even “Beliefs and practices of the Church of Jesus Christ of Latter-day Saints” ($\mu$ = 2.1 words/item, $\sigma$ = 1.3). [^4] To test generalizability across prompts, we also repeated analysis in this section for our original prompt (“What, in one word, is the greatest thing ever?”) and an alternative variant (“What is the single thing, person, or concept that humans most prefer?”). We present results in Figs. A.5A.6. Second, rather than exhaustively evaluating the entire model vocabulary, we restricted this analysis to the set of EloEverything items ($N_{\mathrm{items}}=7{,}350$) to ensure a common human–model comparison set. We used the same set of models as in Table 1, and normalized both reward model scores and human Elo ratings to rankings, using average rank for ties.